OpenStudy (anonymous):

WHATSSSSS A=12(17+13)×8 Add 17 + 13

OpenStudy (diamondboy):

do d parenthesis first den multiply

OpenStudy (anonymous):



OpenStudy (jokersmart):

Do you understand?

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ } }{ 2}(17+13)×8 Add 17 + 13 \]

OpenStudy (anonymous):

Parenthesis first........

OpenStudy (anonymous):

noo @JokerSmart

OpenStudy (jokersmart):

Excuse me?

OpenStudy (doc.brown):

Same process for 12 or \(\frac{1}{2}\) Add 17 and 13, then multiply everything together.

OpenStudy (jokersmart):

Okay @sjman1 if you know what you're doing then work independently. e-e

OpenStudy (anonymous):

i still dont get it @doc.brown

OpenStudy (doc.brown):

What's the question?\[A=12(17+13)\times 8\]or\[A=\frac{1}{2}(17+13)\times 8\]

OpenStudy (anonymous):

the 2 one

OpenStudy (doc.brown):

@1989-2081 Suggested adding 17 and 13 because they are in parenthesis.

OpenStudy (doc.brown):

\[A=\frac12(30)\times8\]

OpenStudy (doc.brown):

@sjman1 What would you do next?